A) What is the mean value of the sum of three independent dice throws?

B) What is the variance of the sum of two independent dice throws?
Answer #1

A) What is the mean value of the sum of three independent dice throws?

1 possible way three dice can total 3
3 ways for 4
6 for 5
10 for 6
15 for 7
21 for 8
25 for 9
27 for 10
27 for 11
25 for 12
21 for 13
15 for 14
10 for 15
6 for 16
3 for 17
1 for 18

Thus, we get

sum 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18
count 1 3 6 10 15 21 25 27 27 25 21 15 10 6 3 1

Mean = (sum*count) for each row / total count

= (3*1 + 4*3 + ... + 18*1) / (1+3+6+..+1)

solving we get

Mean = 2268/216

= 10.5

B) What is the variance of the sum of two independent dice throws?

- Rolling one dice has variance of (35/12)

For two independent dice throws variance = 2*(35/12)

= 5.83